# ObjectClear: Complete Object Removal via Object-Effect Attention

### v1

@@ B More Details of OBER Dataset
Table 4: Comparison of OBER Dataset with Existing Datasets. * indicates that the Counterfactual dataset is not publicly available.
Description RORD [29] MULAN [32] DESOBAv2 [18] Counterfactual [34] OBER (ours)
@@
Effect Mask × × ×
Object-Effect Mask ✓(Coarse) × ×
RGBA Objects × × × ×
@@
Effect Types Shadow × Shadow Shadow Shadow & Reflection
@@
Overview. Table 3 provides an overview of our OBER dataset. The training set consists of 2,715 camera-captured image pairs and 10,000 simulated pairs, all annotated with object masks and object-effect masks. For evaluation, we provide two test subsets: OBER-Test (includes 160 pairs with ground truth) and OBER-Wild (includes 300 pairs without ground truth), where object masks and object-effect masks are also available. In addition, we showcase some samples from our OBER dataset in Fig. 9, including camera-captured data and simulated data, alongside with their annotations such as object masks, object-effect masks, and RGBA object foregrounds.
Reflection Pair Simulation. Thanks to the strong priors of generative models, we observed that the model trained without reflection data was still able to remove some simple reflection effects, a finding consistent with the conclusions drawn in ObjectDrop [34]. However, the model often struggled with more challenging reflection removal cases, such as when the object and its reflection were spatially separated or when the reflection was heavily distorted by surface ripples. To address this limitation, we adopted a human-in-the-loop strategy to collect paired reflection data. Specifically, we used our trained model to perform inference on 200 real-world reflection images and manually selected 50 high-quality results (as shown in Fig. 9(b)), which were then added as an important supplement to the training data. We found that even a small amount of high-quality reflection data pairs could considerably improve the model’s ability to generalize across diverse reflection scenarios.
Comparison with Existing Datasets. We compare our OBER dataset with existing datasets, including those focused on object removal (RORD [29], MULAN [32], Counterfactual Dataset [34]) and one on shadow removal (DESOBA-v2 [18]). As summarized in Table 4, our OBER is the only dataset that includes all three types of mask annotations and RGBA foreground objects, enabling a wide range of tasks such as object removal, effect removal, and joint object-effect removal.
While RORD [29] is among the first to provide object-effect masks, those masks are coarse without separate annotations for objects and their effects. As a result, it cannot support independent removal of either the object or the effect, but only their joint removal. In contrast, our OBER dataset provides precise and separate masks for objects and their effects, enabling more flexible and fine-grained removal tasks. In addition, MULAN [32] focuses solely on object removal without addressing associated effects such as shadows or reflections. DESOBA-v2 [18] targets only effect removal and thus does not handle the objects. Moreover, the ground truth images in both datasets are not captured by cameras but are instead synthesized using generative models, which may limit their realism. The Counterfactual Dataset proposed in ObjectDrop [34] is not publicly available, and it only includes object masks without object-effect masks. It also lacks foreground RGBA data for objects and associated effects, limiting its extensibility and applicability to more advanced object editing tasks.
Figure 9: Samples from our OBER Dataset. We provide some samples from our OBER dataset including both camera-captured data and simulated data. While all data from both categories are annotated with fine-grained object masks and object-effect masks, we also extract RGBA foreground objects with associated effects from camera-captured data (as discussed in Sec. 3.1 in the main paper), which could be used to construct realistic simulated samples. In the simulated samples, we not only include the shadow effects but also the reflection effects, which are rarely included in other datasets (Table 4), thanks to our reflection pair simulation strategy detailed in Sec. B.
C More Results
C.1 Results for Ablation Study
@@
Effectiveness of $\mathcal{L}_{mask}$. Thanks to the rich annotations in our OBER dataset, the object-effect mask enables a supervision loss, denoted as $\mathcal{L}_{mask}$, which guides the cross-attention layers to focus on both the object and its associated effects, while preserving background textures. This supervision facilitates a decoupled optimization of object removal and background reconstruction. As shown in Fig. 10, with the mask loss $\mathcal{L}_{mask}$, the network can adaptively identify the object-effect regions to be removed, as reflected in the attention mask (shown in yellow boxes). This leads to more accurate and complete removal of objects and their shadow effects, without mistakenly erasing unrelated background content.
@@
Figure 10: Effectiveness of $\mathcal{L}_{mask}$. It could be observed that when training without the $\mathcal{L}_{mask}$ supervision, the model struggles to accurately remove the target object and its effects, leading to mistakenly erasing unrelated background object (right) or effects (left). In contrast, when supervised with $\mathcal{L}_{mask}$, the attention mask (shown in yellow boxes) could identify the removal regions well, thus leading to more accurate and complete removal results.
Effectiveness of Simulated Data. To balance data realism and scalability, in addition to the camera-captured data, we scale up our OBER dataset with a carefully designed simulation pipeline. Our simulation data is generated by compositing the foreground RGBA object (extracted from camera-captured data) onto diverse backgrounds. In particular, the simulation of multi-object compositions leads to notable improvements in object removal robustness when mutually occluding objects exist, as shown in Fig. 11 (left). Furthermore, our simulated reflection image pairs (discussed in Sec B) greatly improve the model capability of removing reflections, including challenging cases with significant distortion due to water surface ripples, as shown in Fig. 11 (right).
@@
Figure 11: Effectiveness of Simulated Data. Since our simulation data includes multi-object compositions, training with such data enables the model to accurately remove the target object and its associated effects while preserving unrelated object effects (left). In addition, adding the reflection data pairs during training greatly enhances the model capability of removing reflections, including challenging cases with significant distortion due to water surface ripples (right).
@@
Figure 12: Effectiveness of Attention-Guided Fusion (AG Fusion). We visualize the background detail changes by showing the difference maps between the generated image and the corresponding input. In ideal cases, the difference maps should show large values only in the object removal regions, including the object and its effects, while showing little difference in the unrelated background. It could be observed that when AG Fusion is not employed, although the difference in object removal regions is significant, the difference in the unrelated background is also noticeable. However, with the AG Fusion strategy, we almost eliminate the undesirable background difference thanks to the accurate attention map predicted.
C.2 User Study
@@
To enable a more comprehensive evaluation, we conducted a user study on object removal results for in-the-wild images. We compared ObjectClear with two representative state-of-the-art methods: PowerPaint [43] and RORem [15]. For a fair comparison, we evaluated each baseline method under two settings: (1) conditioned on the object mask and (2) conditioned on the object-effect mask. In contrast, all ObjectClear results were generated using only the object mask, and we compared our outputs against both settings of the baselines.
@@
We invite a total of 30 participants for this user study. Each volunteer was presented 40 randomly selected image quadruples, consisting of: an input image, two results from a baseline method under different mask conditions, and our result. Participants were asked to select the best removal result based on two criteria: the realism of the object region and the preservation of background details.
As summarized in Table 5, ObjectClear outperforms the baselines under both mask settings. Notably, although baseline methods benefited from access to object-effect masks, our ObjectClear still won more user preference with the object mask only.
Table 5: User Study. Our ObjectClear is preferred by human voters over other two representative state-of-the-art methods (PowerPaint [43] and RORem [15]) under both settings. Although the compared methods won more user preference given the object-effect mask instead of the object mask, our ObjectClear still received far more user preference with the object mask only.
Methods PowerPaint (Object Mask) [43] PowerPaint (Object+Effect Mask) [43] ObjectClear (Ours)
User Preference 2.68% 7.03% 90.29%
Methods RORem (Object Mask) [15] RORem (Object+Effect Mask) [15] ObjectClear (Ours)
User Preference 3.51% 11.80% 84.69%
C.3 Results with User Strokes
@@
In practical applications, users often interact with visual systems through imprecise or casually drawn inputs, such as rough scribbles or incomplete masks. These inputs may vary significantly in shape, location, and accuracy. Therefore, it is essential for a robust object removal network to effectively process such arbitrary mask inputs without relying on carefully crafted annotations. Benefiting from our mask augmentation strategy and object-effect attention mechanism, our network demonstrates strong robustness to diverse mask inputs. In this subsection, we simulate user strokes and feed them into the network along with the images. The resulting outputs and attention maps show that our network can accurately identify and attend to the object and its associated effects, even with imprecise masks, as illustrated in Fig. 13.
@@
Figure 13: Results with User Strokes. We simulate user strokes and feed them into the network along with the images. The resulting outputs and attention maps show that our network can accurately identify and attend to the object and its associated effects, even with imprecise masks.
C.4 More Comparisons on In-the-wild Data
In this subsection, we compare ObjectClear with state-of-the-art methods, including SDXL-INP [1], PowerPaint [43], Attentive Eraser [31], DesignEdit [10], and RORem [15], on in-the-wild data. To ensure a fair comparison, we consider two settings: (1) All methods use the same object mask; (2) While other methods use finely annotated object-effect masks, our ObjectClear still uses the object mask. The results under these two settings are shown in Fig. 14 and Fig. 15 separately. Under the first setting, when given the object mask only, our ObjectClear effectively removes shadow and reflection associated with the target object, while all the other methods fail to remove those effects and tend to hallucinate undesirable content or create artifacts. Under the second setting, while other methods do remove the shadow thanks to the given effect region, they also alter or remove the original background content in an undesirable way. In contrast, our ObjectClear removes the object effect accurately with the object mask only while preserving the background with high fidelity.
Figure 14: Object Removal on OBER-Wild - Condition on Object Mask. Given the object mask only, our ObjectClear effectively removes shadow and reflection associated with the target object, while all the other methods fail to remove those effects and tend to hallucinate undesirable content or create artifacts.

### v2

@@ 7 More Details of OBER Dataset
Overview. Table 3 provides an overview of our OBER dataset [link]. The training set consists of 2,715 camera-captured image pairs and 10,000 simulated pairs, all annotated with object masks and object-effect masks. For evaluation, we provide two test subsets: OBER-Test (includes 163 pairs with ground truth) and OBER-Wild (includes 302 pairs without ground truth), where object masks and object-effect masks are also available. Figure 8 presents samples from our OBER dataset, including camera-captured and simulated data with annotations such as object masks, object-effect masks, and RGBA foregrounds.
Reflection Pair Simulation. Thanks to the strong priors of generative models, we observed that a model trained with only limited indoor mirror-reflection data can still generalize to removing simple outdoor reflection effects (e.g., water reflections), which is consistent with the findings in ObjectDrop [38]. However, the model often fails in more challenging cases, particularly when the object and its reflection are spatially separated or when the reflection is heavily distorted by water ripples. To address this limitation, we adopt a human-in-the-loop strategy to collect paired reflection data. Specifically, we first use our trained model to perform inference on 200 real-world reflection images, and then manually select 50 high-quality results (see examples in Fig. 8(b)), which are then added as an important supplement to the training data. We found that even a small amount of high-quality reflection pairs can significantly improve the model’s generalization ability across diverse reflection scenarios.
Comparison with Existing Datasets. We compare our OBER dataset with existing datasets, including those focused on shadow removal (DESOBA-v2 [20]) and object removal (RORD [31], MULAN [35], Counterfactual Dataset [38], Video4Removal [37]). As summarized in Table 4, unlike prior datasets, OBER provides all three types of mask annotations as well as RGBA object foregrounds, enabling a wide range of tasks, including object removal, effect removal, and joint object–effect removal.
DESOBA-v2 [20] targets only effect removal and thus does not handle the objects. In contrast, MULAN [35] focuses solely on object removal without associated effects such as shadows or reflections. Moreover, the ground-truth images in both DESOBA-v2 and MULAN are produced by existing inpainting models, which may limit their visual realism. RORD [31] is among the first to provide object–effect masks, however, these masks are very coarse and can not separately annotate objects and their effects. Consequently, it cannot support independent removal of either the object or the effect. The Counterfactual Dataset [38] and Video4Removal Dataset [37] are designed for object–effect removal tasks, the same with ours, but they are not publicly accessible. Furthermore, these datasets provide only object masks, lacking effect masks and RGBA foregrounds, which limits their scalability and diverse usage. In contrast, our OBER dataset provides richer annotations, including precise and separate masks for objects and their effects, together with RGBA foregrounds, enabling more flexible and fine-grained removal tasks. Notably, the effect masks serve as crucial supervision for learning accurate object–effect removal while preserving background fidelity. OBER is a hybrid dataset comprising high-quality captured and realistic synthetic data, covering diverse and complex scenarios such as multi-object occlusions and indoor/outdoor reflections. We will release our OBER dataset publicly, which we believe will significantly benefit future research in this field.
Table 4: Comparison of OBER Dataset with Existing Datasets. * indicates that the dataset is not publicly available.
Description RORD [31] MULAN [35] DESOBAv2 [20] Counterfactual [38] Video4Removal [37] OBER (ours)
@@
Effect Mask × × × ×
Object-Effect Mask ✓(Coarse) × × ×
RGBA Objects × × × × ×
@@
8 More Results
8.1 Results for Ablation Study
@@
Effectiveness of Adaptive Target-Aware Attention. Thanks to the OBER dataset’s rich annotations, Adaptive Target-Aware Attention (ATA) leverages the object-effect mask and its supervision loss $\mathcal{L}_{mask}$. It guides cross-attention layers to focus on the object and its associated effects while preserving background textures, which enables decoupled optimization of object removal and background reconstruction. As shown in Fig. 9, ATA adaptively identifies object-effect regions to be removed, as reflected in the attention mask (shown in yellow boxes). This leads to more accurate and complete removal of objects and their shadow effects, without mistakenly erasing unrelated background content.
@@
Figure 9: Effectiveness of Adaptive Target-Aware Attention (ATA). It could be observed that when training without ATA (w/o $\mathcal{L}_{mask}$), the model struggles to accurately remove the target object and its effects, leading to mistakenly erasing unrelated background object (right) or effects (left). In contrast, when introducing TAT trained with the mask supervision $\mathcal{L}_{mask}$, the attention maps (shown in yellow boxes) can accurately localize the removal regions, leading to more precise and complete removal results.
@@
Figure 10: Effectiveness of Simulated Data. Since our simulation data includes multi-object compositions, training with such data enables the model to accurately remove the target object and its associated effects while preserving unrelated object effects (left). In addition, adding the reflection data pairs during training greatly enhances the model capability of removing reflections, even in challenging cases (right).
Figure 11: Effectiveness of Spatially-Varying Denoising Strength (SVDS). $\mathrm{DS}=0.99$ often leads to incomplete removal or hallucinated objects, while $\mathrm{DS}=1.0$ causes noticeable color inconsistency (shown after AGF, where the background is from the input image and the object/affected areas are from the removal results). In contrast, SVDS achieves complete object removal with consistent background colors.
Figure 12: Effectiveness of Attention-Guided Fusion (AG Fusion). We visualize background detail changes by computing difference maps between the generated image and input. Ideally, these maps should exhibit large values only within the object removal regions, covering the object and its effects, while remaining minimal in the unrelated background areas. As observed, without AG Fusion, noticeable differences appear in the background. In contrast, with the AG Fusion strategy, these undesired background discrepancies are largely eliminated.
Effectiveness of Simulated Data. To balance realism and scalability, in addition to camera-captured data, we augment OBER with a simulation pipeline. Our simulation data is generated by compositing the RGBA foreground objects (extracted from camera-captured data) onto diverse backgrounds. In particular, the simulation of multi-object compositions leads to notable improvements in object removal robustness under mutual occlusions, as shown in Fig. 10 (left). Furthermore, our simulated reflection pairs greatly improve the model capability of removing reflections, even in challenging cases (Fig. 10, right).
@@
Effectiveness of Spatially-Varying Denoising Strength. In diffusion-based image editing, the initial latent for denoising is typically obtained by adding noise to the input image latent. The denoising strength (DS), $\mathrm{DS}\in[0,1]$, controls the noise level: a larger value injects more noise, thereby pushing the initial noisy latent closer to the pure-noise prior. When $\mathrm{DS}=1.0$, the diffusion process starts entirely from noise, discarding information from the input. In this paper, we propose Spatial-Varying Denoising Strength (SVDS), which applies $\mathrm{DS}=1.0$ within the masked object region and $\mathrm{DS}=0.99$ (an empirical setting commonly adopted by previous methods [26]) outside in the unmasked background, ensuring complete object removal while maintaining color consistency. As shown in Fig. 11, setting $\mathrm{DS}=0.99$ often leads to incomplete removal or hallucinated objects, whereas $\mathrm{DS}=1.0$ results in noticeable color inconsistency. To highlight this inconsistency, the results of $\mathrm{DS}=1.0$ in Fig. 11 are obtained after the Attention-Guided Fusion (AGF) operation, where the background is taken from the original input image, while the object and affected areas are taken from the removal results. In contrast, our method with SVDS achieves superior performance in both object removal and preservation of background color consistency.
Table 5: Quantitative comparison on OmniPaint-Bench and OmniEraser-Bench. The best and second performances are marked in red and orange, respectively.
Methods OmniPaint-Bench OmniEraser-Bench
[…]
SDXL-INP [26] 20.51 0.1924 0.0929 0.1100 21.26 0.1968 0.0938 0.0926
PowerPaint [48] 21.55 0.1882 0.0874 0.0990 22.21 0.2001 0.0911 0.0907
BrushNet [12] 19.00 0.2733 0.1909 0.1446 20.58 0.2098 0.1182 0.1098
DesignEdit [10] 24.32 0.1763 0.0769 0.0717 23.87 0.2168 0.1020 0.0755
CLIPAway [6] 20.07 0.1985 0.0929 0.1129 20.78 0.2035 0.0934 0.0956
FreeCompose [4] 22.25 0.1631 0.0723 0.0770 22.60 0.1782 0.0733 0.0688
Attentive Eraser [34] 24.17 0.1523 0.0501 0.0590 24.77 0.1538 0.0463 0.0388
RORem [18] 23.92 0.1462 0.0478 0.0576 23.70 0.1746 0.0532 0.0416
OmniEraser [37] 23.02 0.2034 0.0564 0.0640 23.83 0.1766 0.0460 0.0481
OmniPaint [42] 25.56 0.1001 0.0249 0.0367 25.67 0.1069 0.0213 0.0182
ObjectClear (Ours) 26.35 0.0950 0.0244 0.0334 27.90 0.0942 0.0230 0.0142
Table 6: Quantitative comparison on MULAN Dataset. The best and second performances are marked in red and orange, respectively.
[…]
SDXL-INP [26] 19.91 0.2494 0.1324 0.1312
PowerPaint [48] 21.18 0.2449 0.1087 0.0962
BrushNet [12] 18.22 0.3181 0.2062 0.1893
DesignEdit [10] 23.26 0.2375 0.1114 0.0725
CLIPAway [6] 20.08 0.2666 0.1180 0.1152
FreeCompose [4] 21.30 0.2337 0.0828 0.0703
Attentive Eraser [34] 23.96 0.1960 0.0551 0.0397
RORem [18] 23.53 0.2369 0.0571 0.0438
OmniEraser [37] 21.56 0.2642 0.0728 0.0682
OmniPaint [42] 22.29 0.1915 0.0650 0.0550
ObjectClear (Ours) 24.89 0.1586 0.0468 0.0373
8.2 Comparisons on Additional Benchmarks
To further evaluate the robustness and generalization ability of our method, we conduct additional evaluations on three benchmarks, i.e., OmniEraser-Bench [37], OmniPaint-Bench [42], MULAN [35].
OmniEraser-Bench and OmniPaint-Bench Datasets. OmniEraser-Bench [37] and OmniPaint-Bench [42] are two recent benchmarks for object–effect removal, which align with our task setting. For a fair comparison, all methods use their default input sizes (OmniEraser [37] at 1024, all others at 512), then resize the outputs to the same size (short side 512) for evaluation. Table 5 shows our approach outperforms all baselines across metrics on both benchmarks.
MULAN Dataset. Since the ground truth of some samples in MULAN [35] retains shadows or reflections, it is not suitable for evaluating object–effect removal. Therefore, we randomly sample 500 “effect-free” image pairs from MULAN for our evaluation. As shown in Table 6, our method achieves the best performance across all metrics, even surpassing RORem [18], which is trained on MULAN, demonstrating the strong object removal capability of our approach. This also confirm that our model performs robustly on “effect-free” object removal datasets, without introducing negative effects when the objects do not show any shadows or reflections.
8.3 Fair Comparison with Object-Effect Mask
Our Attention-Guided Fusion (AGF) module leverages object-effect masks predicted by the proposed Adaptive Target-Aware Attention to blend the original input background back into the generated result. Importantly, these masks are predicted by our model rather than taken from annotations, ensuring that we do not use any privileged information unavailable to other approaches. Furthermore, existing baseline methods do not have the capability to predict object–effect masks and therefore cannot perform background blending in the same way. This makes AGF an integral part of our model design rather than an external post-processing step, and the comparisons in the main paper are therefore fair.
To further demonstrate that the performance gain is not solely due to the background blending, we perform an additional experiment where all baseline methods are given the ground-truth object-effect mask for blending. As shown in Table 7, our method continues to outperform all baselines under this setting, indicating that our superior results primarily come from more effective object–effect generation rather than the availability of blending masks.
Table 7: Quantitative results on RORD-Val with object-effect masks for blending. All baseline methods are equipped with background blending using ground-truth object-effect masks. The best and second performances are marked in red and orange, respectively. Our method using our predicted object-effect masks achieves the best performance across all metrics.
[…]
SDXL-INP [26] (w/b) 21.67 0.1592 0.0688 0.0932
PowerPaint [48] (w/b) 22.51 0.1472 0.0560 0.0606
BrushNet [12] (w/b) 18.48 0.2421 0.1572 0.1465
DesignEdit [10] (w/b) 23.73 0.1580 0.0721 0.0755
CLIPAway [6] (w/b) 21.96 0.1735 0.0666 0.0734
FreeCompose [4] (w/b) 23.08 0.1603 0.0829 0.0834
Attentive Eraser [34] (w/b) 22.90 0.1700 0.0880 0.0854
RORem [18] (w/b) 25.24 0.1398 0.0387 0.0498
ObjectClear (Ours) 26.24 0.1157 0.0191 0.0299
8.4 Generalization to Multi-Object Removal
The OBER dataset also covers multi-object removal cases, and our model generalizes well to such scenarios (Fig. 13). This reflects the strong generalization capability of a network trained on our dataset.
Figure 13: Multi-object removal. Our method removes one or multiple objects simultaneously using masks to indicate the targets.
Figure 14: User Study. Our ObjectClear is preferred by human voters over four representative state-of-the-art methods (PowerPaint [48], RORem [18], OmniEraser [37], and OmniPaint [42].
8.5 User Study
@@
To enable a more comprehensive evaluation, we conducted a user study on object removal results for in-the-wild images. We compared ObjectClear with four representative state-of-the-art methods: PowerPaint [48], RORem [18], OmniEraser [37], and OmniPaint [42]. For a fair comparison, we evaluated PowerPaint and RORem under two settings: (1) conditioned on the object mask and (2) conditioned on the object-effect mask, OmniEraser and OmniPaint are conditioned only on the object mask. All ObjectClear results were generated using only the object mask, and we compared our outputs against both settings of PowerPaint and RORem, as well as the object-mask setting of OmniEraser and OmniPaint.
@@
We invite a total of 30 participants for this user study. Each volunteer was presented 80 randomly selected image quadruples, consisting of: an input image, two results from a baseline method under different mask conditions, and our result (for OmniEraser and OmniPaint, only the object-mask result was provided, consistent with our setting, thus forming a triple set). Participants were asked to select the best removal result based on two criteria: the realism of the object region and the preservation of background details. As summarized in Fig 14, ObjectClear outperforms the baselines under both mask settings. Notably, although some baseline methods benefited from access to object-effect masks, our ObjectClear won more user preference with the object mask only.
8.6 Results with User Strokes
@@
In practical applications, users often interact with visual systems through imprecise or casually drawn inputs, such as rough scribbles or incomplete masks. These inputs may vary significantly in shape, location, and accuracy. Therefore, it is essential for a robust object removal network to effectively process such arbitrary mask inputs without relying on carefully crafted annotations. Benefiting from our mask augmentation strategy and Adaptive Target-Aware Attention mechanism, our network demonstrates strong robustness to diverse mask inputs. In this subsection, we simulate user strokes and feed them into the network along with the images. The resulting outputs and attention maps show that our network can accurately identify and attend to the object and its associated effects, even with imprecise masks, as illustrated in Fig. 15.
@@
Figure 15: Results with User Strokes. We simulate user strokes and feed them into the network along with the images. The resulting outputs and attention maps show that our network can accurately identify and attend to the object and its effects, even with imprecise masks.
8.7 Object Insertion and Movement.
As shown in Fig. 16, even when only the target objects are specified for insertion and movement, ObjectClear is capable of generating plausible and natural shadows and reflections accordingly.
8.8 More Comparisons on In-the-wild Data
In this subsection, we conduct comprehensive comparisons of ObjectClear with state-of-the-art methods across two categories: image inpainting methods and object removal methods, on the in-the-wild data.